OpenStudy (anonymous):

Help me with this one as well! Solve the following simultaneous equation using the substitution method: 3x+8y=24 x+y=3 pleaseeeeeeeeeeeeeeeeee

jhonyy9 (jhonyy9):

x=3-y from second equation and this you need to substitut inside first equation

OpenStudy (anonymous):

Can you do the whole thing? for god's sake?

jhonyy9 (jhonyy9):

3(3-y)+8y=24 9-3y+8y=24 9+5y=24 5y=15 y=15/5 y=3 x=3-y x=3-3 x=0 x=0 y=3